Can someone help me solve and check the following equation?
5x + 8 + 3x - x + 5 = 6x -3
Thanks,
Sher

Answer by funmath(2873)

Hi Sher, is there a Lisa in your class? Because she just asked this question also.
5x+8+3x-x+5=6x-3
5x+3x-x+8+5=6x-3
(5+3-1)x+(8+5)=6x-3
7x+13=6x-3
-6x+7x+13-6x-6x-3
x+13=-3
x+13-13=-3-13
x=-16
----------------
Check: Substitute -16 in for the x's and see if both sides are equal if x=-16.
5(-16)+8+3(-16)-(-16)+5=6(-16)-3
-80+8-48+16+5=-96-3
-99=-99
We're right!!!
Happy calculating!!!!